# CAME-AB: Cross-Modality Attention with Mixture-of-Experts for Antibody Binding Site Prediction

### v1

@@ Abstract
Antibody binding site prediction plays a pivotal role in computational immunology and therapeutic antibody design. Existing sequence or structure methods rely on single-view features and fail to identify antibody-specific binding sites on the antigens—a dual limitation in representation and prediction. In this paper, we propose CAME-AB, a novel Cross-modality Attention framework with a Mixture-of-Experts (MoE) backbone for robust antibody binding site prediction. CAME-AB integrates five biologically grounded modalities—including raw amino acid encodings, BLOSUM substitution profiles, pretrained language model embeddings, structure-aware features, and GCN-refined biochemical graphs—into a unified multimodal representation. To enhance adaptive cross-modal reasoning, we propose an adaptive modality fusion module that learns to dynamically weight each modality based on its global relevance and input-specific contribution. A Transformer encoder combined with an MoE module further promotes feature specialization and capacity expansion. We additionally incorporate a supervised contrastive learning objective to explicitly shape the latent space geometry, encouraging intra-class compactness and inter-class separability. To improve optimization stability and generalization, we apply stochastic weight averaging during training. Extensive experiments on benchmark antibody-antigen datasets demonstrate that CAME-AB consistently outperforms strong baselines on multiple metrics, including Precision, Recall, F1-score, AUC-ROC, and MCC. Ablation studies further validate the effectiveness of each architectural component and the benefit of multimodal feature integration. The model implementation details and the codes are available on https://anonymous.4open.science/r/CAME-AB-C525
@@ 1 Introduction
Predicting antibody binding site—the regions on antigens recognized by specific antibodies—is a critical task in immunological research, vaccine development, and antibody-based therapeutic design ([1]). Accurate identification of these binding sites enables a deeper understanding of immune recognition mechanisms, significantly facilitating rational vaccine and therapeutic antibody engineering ([2, 3]). Traditional approaches predominantly rely on experimentally determined three-dimensional structures or computational sequence-based methods ([4]). However, these methods often face limitations due to structural data scarcity and the inability of sequence-based predictors to capture complex structural relationships essential for accurate predictions ([5]).
@@
Antibody binding site prediction methods based solely on sequence alignment often fail to account for the structural interactions essential to antibody-antigen specificity. For example, sequence-based tools such as BepiPred ([6]), BLAST ([7]) and ClustalW ([8]) are inherently limited in their ability to capture the spatial arrangements and residue interactions critical for accurate prediction. In contrast, structure-based approaches, such as docking simulations and homology modeling, depend heavily on experimentally resolved structures obtained through techniques such as X-ray crystallography or cryo-electron microscopy ([9]). However, these methods are both time-intensive and costly, significantly limiting their widespread applicability.
@@
Recent research in computational biology has shifted attention towards integrating multiple data modalities to address these shortcomings ([10]). Deep learning methods, particularly transformer architectures and graph neural networks (GNNs), have demonstrated exceptional capability in capturing complex, long-range dependencies in both sequential and spatial data ([11]). Transformers have shown great potential in language modeling tasks by effectively modeling long-range dependencies through self-attention mechanisms ([12]). Likewise, GNNs have demonstrated strong capabilities in capturing spatial dependencies not only within protein tertiary structures but also in modeling the topological relationships in protein–protein interaction ([12, 13]), thereby showing the importance of incorporating both structural and interaction-based information in predictive tasks ([14, 13]). The introduction of AlphaFold2 has marked a paradigm shift in protein structure prediction, enabling high-accuracy structural information to be inferred solely from amino acid sequences ([15]). The availability of reliable predicted structures provides an unprecedented opportunity to integrate structural context without requiring experimental resolution, significantly expanding the practical applicability of computational predictions. Additionally, protein language models, such as ProteinBERT ([16]) and evolutionary scale modeling (ESM) ([17]), have demonstrated impressive capability in encoding evolutionary and functional information into dense continuous embeddings. These pretrained models have captured deep evolutionary relationships, contributing significantly to the performance of downstream prediction tasks. Integrating these pretrained embeddings into epitope prediction tasks offers a promising strategy to enhance prediction accuracy and generalizability.
@@
Motivated by recent advances in protein modeling and multimodal learning, we propose CAME-AB, a novel Cross-Modality Attention framework equipped with a Mixture-of-Experts (MoE) backbone for antibody binding site prediction. CAME-AB systematically integrates complementary biological information from multiple representation spaces (as shown in Figure 1). Given the critical role of the CDR region in antigen-antibody recognition ([18]), CAME-AB focuses on the antibody heavy-chain variable region (VH), including CDR-H1, CDR-H2, and CDR-H3 loops, and extracts semantically rich features using diverse encoding strategies. Specifically, CAME-AB incorporates five biologically grounded modalities: (i) one-hot encoding and BLOSUM matrices to capture residue identity and evolutionary substitution patterns; (ii) pretrained contextual embeddings from a large protein language model, i.e., ESMC; (iii) structural features derived from ESM’s structure-aware output layers; and (iv) residue-level biochemical similarity graphs constructed using PyBioMed descriptors, from which we obtain structural-aware node embeddings via a graph convolutional network. These heterogeneous representations are projected into a unified latent space and fused via a learnable adaptive modality fusion module that jointly models modality informativeness, sample-specific variation, and class-aware semantic priors. To improve discriminative capacity and model generalization, our proposed architecture incorporates three additional components: (1) a Mixture-of-Experts (MoE) module to encourage feature specialization across latent subspaces; (2) a supervised contrastive learning objective to enforce intra-class compactness and inter-class separability in the embedding space; and (3) Stochastic Weight Averaging (SWA) for optimization smoothing and enhanced generalization. We evaluate our framework on multiple public antibody-antigen binding datasets. Extensive experiments show that our method consistently outperforms state-of-the-art baselines on multiple metrics. Ablation studies further verify the effectiveness of each multi-view feature and architectural component.
@@ 2.1 Antibody Binding Site Prediction
  • Sequence-based models (e.g., BepiPred ([6]), ABCPred ([19])) employ machine learning on sequence features to predict linear epitopes. Although computationally efficient, they ignore spatial context and fail to capture conformational epitopes.
  • Structure-based methods (e.g., Molecular Docking ([20])) utilize 3D structural information through docking simulations and geometric analysis. Although better at identifying spatial epitopes, these approaches require experimentally resolved structures that are often unavailable.
@@ 2.2 Feature Representation in Bioinformatics
Feature representation is a cornerstone of bioinformatics, enabling the extraction and integration of meaningful patterns from biological data. In this section, we introduce key feature representation methods, including ESMC, One-hot encoding, BLOSUM ([21]), and PyBioMed ([22]), alongside a discussion of multi-view learning approaches and their relevance to bioinformatics tasks. A detailed description of each feature representation and its bioinformatics relevance is provided in the Supplementary Material (Section 1).
@@ 3.1 Problem Formulation and Multimodal Representation
To address this, we formulate epitope prediction as a multimodal learning task. Our framework integrates five biologically grounded feature modalities as shown in Figure 2: (i) Amino acid encoding schemes: we incorporate one-hot encoding to preserve raw residue identity, BLOSUM substitution matrices to model evolutionary conservation patterns, and contextualized embeddings derived from pretrained protein language models such as ESMC to capture sequence semantics and long-range dependencies; (ii) Structure-informed representations: we utilize the structural output layer of ESMC as a dedicated structural modality. This layer provides an approximate estimation of spatial residue relationships even in the absence of experimentally resolved structures; (iii) Graph-based biochemical features: we construct a residue-level graph where each node corresponds to an amino acid and is initialized using its ESMC sequence embedding. Edges are established between residue pairs based on pairwise biochemical similarity, computed from PyBioMed-derived physicochemical descriptors such as hydrophobicity, polarity, and surface accessibility. This graph is used to train a Graph Convolutional Network (GCN) ([23]), for refining node embeddings by aggregating spatial and chemical context from neighboring residues.
@@ 3.2.1 Multimodal Feature Encoding
For the GCN branch, we construct a residue-level graph $G=(V,E)$, where each node $v_{i}\in V$ corresponds to a residue and is initialized using its ESMC embedding. Edges $(v_{i},v_{j})\in E$ are established based on pairwise biochemical similarity, computed using PyBioMed-derived descriptors such as hydrophobicity, polarity, and charge. residues are connected if their pairwise similarity, computed using selected PyBioMed descriptors, exceeds a threshold. Given the graph $G$, we apply a two-layer GCN ([23]) to refine the node embeddings. Let $\mathbf{X}\in\mathbb{R}^{L\times d}$ be the input node feature matrix and $\mathbf{A}\in\mathbb{R}^{L\times L}$ be the adjacency matrix of the graph $G$, the GCN layer is defined as ([23]):
@@ 3.2.2 Adaptive Modality Fusion
Inspired by hierarchical gating and label-aware conditioning to mitigate spurious modality correlations ([24]), our approach introduces three types of adaptive weights. Let $\tilde{F}^{(m)}\in\mathbb{R}^{L\times d}$ denote the projected embedding of the $m$-th modality. The fused representation $\mathbf{F}_{\text{fused}}\in\mathbb{R}^{L\times d}$ is computed as a weighted sum over all $M$ modalities:
@@ 3.2.3 Transformer-based Representation Encoding with Mixture-of-Experts
After adaptive fusion, the integrated representation $\mathbf{F}_{\text{fused}}\in\mathbb{R}^{L\times d}$ is passed through a two-layer Transformer encoder to capture intra-sequence dependencies and inter-residue interactions across CDRs. We adopt a Pre-LayerNorm (Pre-LN) architecture ([25]) for improved training stability, defined as:
@@
To further enhance feature specialization and model capacity, we introduce a Mixture-of-Experts (MoE) module ([26]). It consists of $K$ expert networks $\{E_{k}\}_{k=1}^{K}$, each implemented as a two-layer MLP. A gating network assigns a soft distribution over experts ([26]):
@@
To prevent expert collapse and encourage diverse specialization, we introduce a diversity regularization loss ([27]):
@@ 3.2.4 Prediction and Contrastive Embedding
Let $P(i)\subseteq\{1,\ldots,N\}$ denote the set of indices of positive samples in the batch that share the same ground-truth label as sample $i$, and let $A(i)=\{1,\ldots,N\}\setminus\{i\}$ be the set of all other samples excluding $i$ itself. The supervised contrastive loss is then defined as ([28]):
@@
To stabilize training, we apply Stochastic Weight Averaging (SWA) ([29]) during the final training phase:
@@
The focal loss function was originally proposed to address class imbalance by down-weighting easy examples and focusing learning on hard, misclassified samples, as defined as ([30]):
@@ 4.1.1 Dataset and Preprocessing
To ensure fair evaluation and reproducibility, we construct dataset following a standardized pipeline in ABS ([31]). Specifically, we collect antibody sequences annotated with antigen-binding information, focusing on the VH, CDR1, CDR2, and CDR3 regions. Sequences containing missing or incomplete entries are discarded to ensure data integrity. To mitigate redundancy and reduce potential data leakage, we perform sequence clustering using CD-HIT ([32]) at a 90% identity threshold. Clusters are further refined using an 80% similarity threshold ([33]) to construct non-overlapping training, validation, and test partitions. This stratification guarantees that highly similar sequences do not appear across different splits, thereby promoting a robust generalization evaluation. To address class imbalance, we apply a combination of noise reduction and controlled up/down-sampling strategies within each split. The resulting dataset is divided into 80% for training, 10% for validation, and 10% for testing. Additionally, we remove all duplicate entries post-clustering to ensure that each sequence instance contributes uniquely to model training and evaluation.
@@ 4.3 Comparison with Baselines
To evaluate the effectiveness of our proposed approach, we compare it against several state-of-the-art (SOTA) methods, including ABS ([31]), ME-ACP ([34]), xDeep-AcPEP ([35]), and PreAlgPro ([36]).
@@
ABS ([31]) 0.8041±0.0024 0.6934±0.0025 0.7336±0.0024 0.9316±0.0004 0.5528±0.0037
ME-ACP ([34]) 0.8122±0.0031 0.8198±0.0026 0.8140±0.0029 0.9587±0.0005 0.7035±0.0044
xDeep-AcPEP ([35]) 0.7944±0.0126 0.8023±0.0118 0.7955±0.0105 0.9093±0.0067 0.6754±0.0180
PreAlgPro ([36]) 0.7754±0.0112 0.7756±0.0071 0.7749±0.0085 0.9332±0.0023 0.6354±0.0125
@@ A.3 BLOSUM: Block Substitution Matrix
The BLOSUM family of matrices, such as BLOSUM62 ([21]), provides a scoring system based on observed substitutions in conserved regions of proteins. These matrices incorporate evolutionary information and are often used in sequence alignment and similarity-based feature extraction, offering insights into residue-level functional importance.

### v2

@@ Abstract
Antibody binding site prediction plays a pivotal role in computational immunology and therapeutic antibody design. Existing sequence or structure methods rely on single-view features and fail to identify antibody-specific binding sites on the antigens. In this paper, we propose CAME-AB, a novel Cross-modality Attention framework with a Mixture-of-Experts (MoE) backbone for robust antibody binding site prediction. CAME-AB integrates five biologically grounded modalities, including raw amino acid encodings, BLOSUM substitution profiles, pretrained language model embeddings, structure-aware features, and GCN-refined biochemical graphs, into a unified multimodal representation. To enhance adaptive cross-modal reasoning, we propose an adaptive modality fusion module that learns to dynamically weight each modality based on its global relevance and input-specific contribution. A Transformer encoder combined with an MoE module further promotes feature specialization and capacity expansion. We additionally incorporate a supervised contrastive learning objective to explicitly shape the latent space geometry, encouraging intra-class compactness and inter-class separability. To improve optimization stability and generalization, we apply stochastic weight averaging during training. Extensive experiments on benchmark antibody-antigen datasets demonstrate that CAME-AB consistently outperforms strong baselines on multiple metrics, including Precision, Recall, F1-score, AUC-ROC, and MCC. Ablation studies further validate the effectiveness of each architectural component and the benefit of multimodal feature integration. The model implementation details and the codes are available on https://anonymous.4open.science/r/CAME-AB-C525
@@ 1 Introduction
Predicting antibody binding site—the regions on antigens recognized by specific antibodies—is a critical task in immunological research, vaccine development, and antibody-based therapeutic design [1]. Accurate identification of these binding sites enables a deeper understanding of immune recognition mechanisms, significantly facilitating rational vaccine and therapeutic antibody engineering [2, 3]. Traditional approaches predominantly rely on experimentally determined three-dimensional structures or computational sequence-based methods [4]. However, these methods often face limitations due to structural data scarcity and the inability of sequence-based predictors to capture complex structural relationships essential for accurate predictions [5].
@@
Antibody binding site prediction methods based solely on sequence alignment often fail to account for the structural interactions essential to antibody-antigen specificity. For example, sequence-based tools such as BepiPred [6], BLAST [7] and ClustalW [8] are inherently limited in their ability to capture the spatial arrangements and residue interactions critical for accurate prediction. In contrast, structure-based approaches, such as docking simulations and homology modeling, depend heavily on experimentally resolved structures obtained through techniques such as X-ray crystallography or cryo-electron microscopy [9]. However, these methods are both time-intensive and costly, significantly limiting their widespread applicability.
@@
Recent research in computational biology has shifted attention towards integrating multiple data modalities to address these shortcomings [10]. Deep learning methods, particularly transformer architectures and graph neural networks (GNNs), have demonstrated exceptional capability in capturing complex, long-range dependencies in both sequential and spatial data [11]. Transformers have shown great potential in language modeling tasks by effectively modeling long-range dependencies through self-attention mechanisms [12]. Likewise, GNNs have demonstrated strong capabilities in capturing spatial dependencies not only within protein tertiary structures but also in modeling the topological relationships in protein–protein interaction [12, 13], thereby showing the importance of incorporating both structural and interaction-based information in predictive tasks [14, 13]. The introduction of AlphaFold2 has marked a paradigm shift in protein structure prediction, enabling high-accuracy structural information to be inferred solely from amino acid sequences [15]. The availability of reliable predicted structures provides an unprecedented opportunity to integrate structural context without requiring experimental resolution, significantly expanding the practical applicability of computational predictions. Additionally, protein language models, such as ProteinBERT [16] and evolutionary scale modeling (ESM) [17], have demonstrated impressive capability in encoding evolutionary and functional information into dense continuous embeddings. These pretrained models have captured deep evolutionary relationships, contributing significantly to the performance of downstream prediction tasks. Integrating these pretrained embeddings into epitope prediction tasks offers a promising strategy to enhance prediction accuracy and generalizability.
@@
Motivated by recent advances in protein modeling and multimodal learning, we propose CAME-AB, a novel Cross-Modality Attention framework equipped with a Mixture-of-Experts (MoE) backbone for antibody binding site prediction. CAME-AB systematically integrates complementary biological information from multiple representation spaces (as shown in Figure 1). Given the critical role of the CDR region in antigen-antibody recognition [18], CAME-AB focuses on the antibody heavy-chain variable region (VH), including CDR-H1, CDR-H2, and CDR-H3 loops, and extracts semantically rich features using diverse encoding strategies. Specifically, CAME-AB incorporates five biologically grounded modalities: (i) one-hot encoding and BLOSUM matrices to capture residue identity and evolutionary substitution patterns; (ii) pretrained contextual embeddings from a large protein language model, i.e., ESMC; (iii) structural features derived from ESM’s structure-aware output layers; and (iv) residue-level biochemical similarity graphs constructed using PyBioMed descriptors, from which we obtain structural-aware node embeddings via a graph convolutional network. These heterogeneous representations are projected into a unified latent space and fused via a learnable adaptive modality fusion module that jointly models modality informativeness, sample-specific variation, and class-aware semantic priors. To improve discriminative capacity and model generalization, our proposed architecture incorporates three additional components: (1) a Mixture-of-Experts (MoE) module to encourage feature specialization across latent subspaces; (2) a supervised contrastive learning objective to enforce intra-class compactness and inter-class separability in the embedding space; and (3) Stochastic Weight Averaging (SWA) for optimization smoothing and enhanced generalization. We evaluate our framework on multiple public antibody-antigen binding datasets. Extensive experiments show that our method consistently outperforms state-of-the-art baselines on multiple metrics. Ablation studies further verify the effectiveness of each multi-view feature and architectural component.
@@ 2.1 Antibody Binding Site Prediction
  • Sequence-based models (e.g., BepiPred [6], ABCPred [19]) employ machine learning on sequence features to predict linear epitopes. Although computationally efficient, they ignore spatial context and fail to capture conformational epitopes.
  • Structure-based methods (e.g., Molecular Docking [20]) utilize 3D structural information through docking simulations and geometric analysis. Although better at identifying spatial epitopes, these approaches require experimentally resolved structures that are often unavailable.
@@ 2.2 Feature Representation in Bioinformatics
Feature representation is a cornerstone of bioinformatics, enabling the extraction and integration of meaningful patterns from biological data. In this section, we introduce key feature representation methods, including ESMC, One-hot encoding, BLOSUM [21], and PyBioMed [22], alongside a discussion of multi-view learning approaches and their relevance to bioinformatics tasks. A detailed description of each feature representation and its bioinformatics relevance is provided in the Supplementary Material (Section A).
@@ 3.1 Problem Formulation and Multimodal Representation
To address this, we formulate epitope prediction as a multimodal learning task. Our framework integrates five biologically grounded feature modalities as shown in Figure 2: (i) Amino acid encoding schemes: we incorporate one-hot encoding to preserve raw residue identity, BLOSUM substitution matrices to model evolutionary conservation patterns, and contextualized embeddings derived from pretrained protein language models such as ESMC to capture sequence semantics and long-range dependencies; (ii) Structure-informed representations: we utilize the structural output layer of ESMC as a dedicated structural modality. This layer provides an approximate estimation of spatial residue relationships even in the absence of experimentally resolved structures; (iii) Graph-based biochemical features: we construct a residue-level graph where each node corresponds to an amino acid and is initialized using its ESMC sequence embedding. Edges are established between residue pairs based on pairwise biochemical similarity, computed from PyBioMed-derived physicochemical descriptors such as hydrophobicity, polarity, and surface accessibility. This graph is used to train a Graph Convolutional Network (GCN) [23], for refining node embeddings by aggregating spatial and chemical context from neighboring residues.
@@ 3.2.1 Multimodal Feature Encoding
For the GCN branch, we construct a residue-level graph $G=(V,E)$, where each node $v_{i}\in V$ corresponds to a residue and is initialized using its ESMC embedding. Edges $(v_{i},v_{j})\in E$ are established based on pairwise biochemical similarity, computed using PyBioMed-derived descriptors such as hydrophobicity, polarity, and charge. residues are connected if their pairwise similarity, computed using selected PyBioMed descriptors, exceeds a threshold. Given the graph $G$, we apply a two-layer GCN [23] to refine the node embeddings. Let $\mathbf{X}\in\mathbb{R}^{L\times d}$ be the input node feature matrix and $\mathbf{A}\in\mathbb{R}^{L\times L}$ be the adjacency matrix of the graph $G$, the GCN layer is defined as [23]:
@@ 3.2.2 Adaptive Modality Fusion
Inspired by hierarchical gating and label-aware conditioning to mitigate spurious modality correlations [24], our approach introduces three types of adaptive weights. Let $\tilde{F}^{(m)}\in\mathbb{R}^{L\times d}$ denote the projected embedding of the $m$-th modality. The fused representation $\mathbf{F}_{\text{fused}}\in\mathbb{R}^{L\times d}$ is computed as a weighted sum over all $M$ modalities:
@@ 3.2.3 Transformer-based Representation Encoding with Mixture-of-Experts
After adaptive fusion, the integrated representation $\mathbf{F}_{\text{fused}}\in\mathbb{R}^{L\times d}$ is passed through a two-layer Transformer encoder to capture intra-sequence dependencies and inter-residue interactions across CDRs. We adopt a Pre-LayerNorm (Pre-LN) architecture [25] for improved training stability, defined as:
@@
To further enhance feature specialization and model capacity, we introduce a Mixture-of-Experts (MoE) module [26]. It consists of $K$ expert networks $\{E_{k}\}_{k=1}^{K}$, each implemented as a two-layer MLP. A gating network assigns a soft distribution over experts [26]:
@@
To prevent expert collapse and encourage diverse specialization, we introduce a diversity regularization loss [27]:
@@ 3.2.4 Prediction and Contrastive Embedding
Let $P(i)\subseteq\{1,\ldots,N\}$ denote the set of indices of positive samples in the batch that share the same ground-truth label as sample $i$, and let $A(i)=\{1,\ldots,N\}\setminus\{i\}$ be the set of all other samples excluding $i$ itself. The supervised contrastive loss is then defined as [28]:
@@
To stabilize training, we apply Stochastic Weight Averaging (SWA) [29] during the final training phase:
@@
The focal loss function was originally proposed to address class imbalance by down-weighting easy examples and focusing learning on hard, misclassified samples, as defined as [30]:
@@ 4.1.1 Dataset and Preprocessing
To ensure fair evaluation and reproducibility, we construct dataset following a standardized pipeline in ABS [31]. Specifically, we collect antibody sequences annotated with antigen-binding information, focusing on the VH, CDR1, CDR2, and CDR3 regions. Sequences containing missing or incomplete entries are discarded to ensure data integrity. To mitigate redundancy and reduce potential data leakage, we perform sequence clustering using CD-HIT [32] at a 90% identity threshold. Clusters are further refined using an 80% similarity threshold [33] to construct non-overlapping training, validation, and test partitions. This stratification guarantees that highly similar sequences do not appear across different splits, thereby promoting a robust generalization evaluation. To address class imbalance, we apply a combination of noise reduction and controlled up/down-sampling strategies within each split. The resulting dataset is divided into 80% for training, 10% for validation, and 10% for testing. Additionally, we remove all duplicate entries post-clustering to ensure that each sequence instance contributes uniquely to model training and evaluation.
@@ 4.3 Comparison with Baselines
To evaluate the effectiveness of our proposed approach, we compare it against several state-of-the-art (SOTA) methods, including ABS [31], ME-ACP [34], xDeep-AcPEP [35], and PreAlgPro [36].
@@
ABS [31] 0.8041±0.0024 0.6934±0.0025 0.7336±0.0024 0.9316±0.0004 0.5528±0.0037
ME-ACP [34] 0.8122±0.0031 0.8198±0.0026 0.8140±0.0029 0.9587±0.0005 0.7035±0.0044
xDeep-AcPEP [35] 0.7944±0.0126 0.8023±0.0118 0.7955±0.0105 0.9093±0.0067 0.6754±0.0180
PreAlgPro [36] 0.7754±0.0112 0.7756±0.0071 0.7749±0.0085 0.9332±0.0023 0.6354±0.0125
@@ A.3 BLOSUM: Block Substitution Matrix
The BLOSUM family of matrices, such as BLOSUM62 [21], provides a scoring system based on observed substitutions in conserved regions of proteins. These matrices incorporate evolutionary information and are often used in sequence alignment and similarity-based feature extraction, offering insights into residue-level functional importance.